# MathCAMPS: Fine-grained Synthesis of Mathematical Problems From Human Curricula

### v1

@@ Abstract
Mathematical problem solving is an important skill for Large Language Models (LLMs), both as an important capability and a proxy for a range of reasoning abilities. Existing benchmarks probe a diverse set of skills, but they yield aggregate accuracy metrics, obscuring specific abilities or weaknesses. Furthermore, they are difficult to extend with new problems, risking data contamination over time. To address these challenges, we propose MathCAMPS: a method to synthesize high-quality mathematical problems at scale, grounded on 44 fine-grained “standards” from the Mathematics Common Core (CC) Standard for K-8 grades. We encode each standard in a formal grammar, allowing us to sample diverse symbolic problems and their answers. We then use LLMs to realize the symbolic problems into word problems. We propose a cycle-consistency method for validating problem faithfulness. Finally, we derive follow-up questions from symbolic structures and convert them into follow-up word problems—a novel task of mathematical dialogue that probes for robustness in understanding. Experiments on 23 LLMs show surprising failures even in the strongest models (in particular when asked simple follow-up questions). Moreover, we evaluate training checkpoints of Pythia 12B on MathCAMPS, allowing us to analyze when particular mathematical skills develop during its training. Our framework enables the community to reproduce and extend our pipeline for a fraction of the typical cost of building new high-quality datasets.
@@ 1 Introduction
As Large Language Models (LLMs) become increasingly capable, mathematical reasoning problems have emerged as a key benchmark for evaluating their abilities. Mathematical reasoning is a critical subproblem of many important tasks, such as scientific question answering and quantitative data analysis, making it a prerequisite for a range of downstream applications. Moreover, mathematical reasoning tests a broad spectrum of reasoning skills, serving as a valuable proxy for assessing reasoning capabilities more generally. Consequently, several benchmarks, notably GSM8K [9] and MATH [14], became popular measures of the progress of LLMs, with each new generation of models demonstrating rapid advancements. Equal contribution.
However, the classical approach to benchmarking in Machine Learning, which involves evaluating models on a fixed set of human-created problems, faces new fundamental challenges in the era of LLMs. First, these models are trained on massive public datasets that may unintentionally include the very benchmarks used for evaluation, raising concerns about data contamination [26, 8, 4]. This problem is exacerbated by the lack of access to the training data of most state-of-the-art LLMs, such as GPT-4 [1], Claude [2], and even open-weight models, such as LLaMA [25]. Evaluating LLMs on novel problems could mitigate the data contamination concerns. But creating new mathematical problems is challenging. Crafting new high-quality problems requires expertise and is expensive; sourcing problems from public sources does not address the question of whether LLMs might have been trained on those problems.
Moreover, while existing benchmarks serve to track overall progress in the field, they do not inform us about what mathematical abilities current language models do and do not have. A single aggregate accuracy — in a topic as diverse as mathematics — does not provide insights into specific capabilities or challenges for current language models, and how those have been changing over time. For instance, GPT-4 [1] improved by 35% on GSM8K when compared to GPT-3 [7]; yet, it is still challenging to understand which improved capabilities might have accounted for this improvement (e.g., arithmetic with larger numbers, proficiency with fractions or decimals, or understanding of longer problems). Such an analysis would help shed light on open questions about language model learning, and how it relates to (or diverges from) human learning.
@@
To address these challenges, we propose the Mathematics Common Core Assessment of Problem Solving — MathCAMPS — a framework for synthesizing high-quality mathematical word problems at scale. Our approach is grounded in the Mathematics Common Core (CC) Standards from Kindergarten through 8th grade. The CC standardizes a mathematics curriculum adopted by thousands of schools, describing specific abilities that students should learn by each grade. By constructing MathCAMPS in direct relation to the CC, our benchmark enables a series of rich analyses of mathematical proficiency in language models, allowing direct parallels to abilities that human students are also evaluated on. We encode the skills described in the CC (namely the standards) in a grammar that allows us to sample an arbitrary number of diverse problems targeting that skills (e.g., word problems involving addition of decimals, or solving systems of equations with fractions), represented symbolically.
Our pipeline uses a symbolic solver (SymPy) to obtain answers to the symbolic problems, and employs an LLM to realize those into word problems. We introduce a cycle-consistency method to validate whether a word problem faithfully represents the original symbolic problem. Prompting the LLM to back-translate the word problem into a symbolic structure and comparing the new answer to the original enables us to eliminate most unfaithful generation errors and maintain high quality.
Furthermore, building on our symbolic representation of problem structures, we introduce a novel task of “mathematical dialogue”. In this task, once the LLM answers a problem correctly, we ask a follow-up question to further probe understanding. We introduce two types of follow-up problems: counterfactual, where we modify an aspect of the original problem and request an updated answer, and incremental, where we provide additional information and ask for a new answer. These questions require simultaneously understanding the original problem and the LLM’s own solution — an additional challenge that several models struggle with.
Using our framework, we synthesize problems for each of 44 CC standards, resulting in a dataset of 4,900 initial problems. We also generate follow-up questions (incremental and conterfactual) in standards where those apply, yielding 9607 total problems. We evaluate a suite of 23 language models, both proprietary and open. Our analysis uncovers surprising failures, particularly in response to simple follow-up questions, revealing notable gaps even in strong models. Moreover, to the best of our knowledge, we perform the first analysis of the learning dynamics of mathematical skills during LLM training, leveraging checkpoints of Pythia 12B [6]. Our contributions are:
  • We present MathCAMPS, a framework for synthesizing high-quality mathematical word problems at scale, stratified into fine-grained capabilities defined by the Mathematics Common Core Standards for K-8 grades. We release 9607 problems and our extensible pipeline to generate arbitrarily many more.
  • We introduce a cycle-consistency method to validate the faithfulness of the generated word problems to their underlying symbolic structures.
  • We propose a novel task of “mathematical dialogue,” featuring counterfactual and incremental follow-up questions that probe the models’ understanding more deeply.
  • We evaluate a diverse set of 23 language models on our dataset, revealing surprising failures and gaps in performance, even in strong models.
@@ 2 Related Work
Our work closely relates to (i) current benchmarks of mathematical reasoning in LLMs, (ii) benchmarks constructed using LLMs, and (iii) behavioral testing and applications in NLP.
@@
MATH [14] and GSM8K [9] have been two leading benchmarks for the evaluation of mathematical reasoning in LLMs. Both datasets consist entirely of human-authored problems — a process that is expensive to reproduce, and as a result, neither benchmarks were updated since their initial releases. Given that LLMs are trained on Web data, it is unclear whether they might have been trained on the test problems of these benchmarks [8] – either directly or from other sources (e.g., all problems in MATH come from past public competitions). In fact, GSM1K [26], a new dataset that independently attempted to reproduce the data distribution of GSM8K, has found reduced performance on several models, suggesting test set contamination.
@@
As collecting data from human annotaotors at scale is expensive (especially in domains requiring expertise, such as mathematics), prior work has relied on LLMs to aid the generation of large-scale benchmarks [12]. BigToM [10], a benchmark of social reasoning in LLMs, applied the idea of symbolically scaffolding questions for the LLM to realize in natural language, an approach that we transport to mathematics. Dyval [27] proposed a method for generating reasoning problems for LLMs based on a DAG representing the computation. While Dyval contains two mathematical tasks (arithmetic and solving linear equations), MathCAMPS takes this idea further for mathematical reasoning, spanning 44 skills directly grounded on a human curriculum.
@@
Our goal to provide a fine-grained evaluation of mathematical reasoning has parallels with behavioral testing — the idea of testing software systems on specific features, as opposed to just their overall adequacy [20]. In particular, CheckList [20] allowed testing machine translation models for fine-grained failure modes. Dynaboard [18] proposed an NLP leaderboard where users can adapt to their own needs by choosing the utility of different metrics; our dataset enables a similar user-customizable comparison between models for mathematical reasoning.
@@ 3.1 The Mathematics Common Core
We take 44 standards spanning grades K through 8 to compose MathCAMPS, focusing on standards that are amenable to automatic problem generation with a final answer in text form. The complete CC curriculum has 229 standards across grades K through 8, bring our coverage to 19.2% of the curriculum for these grades. Notably, we currently do not cover standards focusing on conceptual understanding (e.g., 3.OA.D.9 – “Identify arithmetic patterns […], and explain them using properties of operations.”), or standards that emphasize visual reasoning (e.g., 6.G.A.4 – “Represent three-dimensional figures using nets made up of rectangles and triangles, and use the nets to find the surface area of these figures.”). All 44 standards covered in MathCAMPS are listed in Appendix A.
@@
We represent CC standards as non-terminals in an attribute grammar [13] — a rich formalism that can encode semantic, context-sensitive rules. Attribute grammars can encode syntax much like a context-free grammar, but also allow us to embed information processing (e.g., setting and testing conditions on attributes, such as bounds on constants) in the production rules. We map each standard $s$ to a non-terminal $P_{s}$, such that all strings produced by expanding $P_{s}$ using production rules are valid symbolic representations of a problem pertaining to standard $i$. Figure 1 shows a (simplified) grammar for the standard 1.OA.A.1 – “Use addition and subtraction within 20 to solve word problems involving situations of adding to, taking from, putting together”. Here, a word problem, generated by the Problem non-terminal, consists of a sequence of declarative statements expressing equations between expressions. For this standard, an expression consists of addition, subtraction, variables, and constants. After these declarations, the problem ends with a question — an expression representing the value that the problem asks for. Concretely, our grammar is implemented in Python: each non-terminal becomes a stochastic function that samples and applies a production rule, recursively expanding non-terminals that it produces. In the grammar in Figure 1 (A), sampling a Problem generates a structure such as the one shown in Figure 1 (B).
@@
When sampling problems, there is no a priori guarantee that all generated statements are necessary to answer the question. To avoid such statements, we remove them by applying a simple graph reachability algorithm on a dependency graph between statements, removing statements that the answer does not depend on. This enforces the constraint of only having useful statements in problems. Besides this constraint, which we always enforce, each standard can apply specific constraints. The standard 1.OA.A.1 has an example of such constraint: it requires that students only be asked to use “addition and subtraction within $20$.” To be faithful to this standard, we must validate that no intermediate values used in the solution exceed $20$. To encode this and other constraints across the curriculum, we implement a suite of $6$ parameterized filters (detailed in Appendix C) that are selectively applied depending on the standard’s specification. Applying rejection sampling from the grammar using the standard’s filters gives a procedure for generating valid symbolic problems. For all standards that can be formulated as solving a system of linear equations, we use SymPy [19] to obtain final answers. For other cases, we use two simple custom procedures (to list the factors of numbers and to compare values).
@@ 3.2 From symbolic to word problems
Unlike generating problem stories from a fixed set of templates, using a language model for generating natural language problems gives us fluid, diverse language. Unfortunately, we also lose any guarantee that the generated word problem represents the original symbolic structure faithfully. To mitigate this issue, we also introduce a cycle consistency method that we have found to drastically improve problem quality. Precisely, we use the same few-shot examples we crafted for each standard in reverse (i.e., with the natural language problem coming first, followed by the symbolic structure) to have GPT-4 translate the word problem it wrote into a symbolic structure. In this step, the model is not given the original structure. We then parse and apply the appropriate solver to the generated symbolic problem; we consider the generation cycle-consistent if the answers to the original and recovered problems are the same (illustrated in Figure 1). We then discard problems that fail this test.
This cycle consistency test significantly improves the reliability of our pipeline. We manually evaluated 245 random problems generated by sampling a symbolic structure and then a word problem from GPT-4. Out of those, we identified 30 word problems (12.2%) that were not faithful to the original symbolic structure — for those, the answer that we compute to the symbolic problem does not match our manual solution to the word problem. Cycle consistency discarded 25 of those (and 7 problems that were indeed faithful). Out of the remaining 215 problems, 210 (97.7%) were judged as faithful in our manual check.

### v2

# From Next-Token to Mathematics: The Learning Dynamics of Mathematical Reasoning in Language Models

@@ Abstract
Large Language Models (LLMs) solely trained on next-token prediction learn to solve a wide range of problems involving mathematical reasoning. But how does this ability evolve during training? We show the first analysis of how mathematical reasoning abilities of several open-weight LLMs develop during pre-training and post-training. To this end, we construct MathCAMPS  [^1] All code and data associated with the project can be accessed at https://github.com/gpoesia/mathcamps, a synthetic dataset of novel mathematical reasoning problems grounded in 44 fine-grained skills taken from the Common Core curriculum from K to 8th grades. In one experiment, we show that mathematical skills are learned during pre-training in an order that measurably correlates with the human-designed curriculum, even though training data are randomly ordered. We also show a detailed analysis of which mathematical abilities benefit from instruction tuning, a widely used post-training method and, in contrast, which skills suffer. Our work paves the way for an empirical understanding of LLM training dynamics in relation to reasoning.
@@ 1 Introduction
Large language models are pre-trained on a simple objective of next-token prediction. From this pretext task alone they acquire the ability to perform tasks requiring a surprising range of skills for humans, including answering complex questions, translating between languages, and generating code. Among these tasks, one notable case that has been widely used as a proxy for the general capability of new generations of LLM has been their ability to perform mathematical reasoning. This attention to mathematics is natural, since mathematics is a key tool for reasoning underlying many of the most impressive feats of human engineering, such as sending rockets to space or building nuclear reactors. Many benchmarks, such as GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), or more recently, FrontierMath (Glazer et al., 2024), have been proposed to evaluate LLMs on mathematical problems of various levels. But how this ability develops during training is still poorly understood.
How can we gain insight into the learning dynamics of mathematical reasoning skills? Several open-weights LLMs have been made available with intermediate checkpoints, including Pythia (Biderman et al., 2023), OLMo (OLMo et al., 2025) and Amber (Liu et al., 2023). However, several challenges arise in designing informative evaluations of these models during their training. First, we would like to rule out the potential effect of data contamination: LLMs being able to answer a question simply because it was seen during pre-training. Unfortunately, since several of the well-known mathematical reasoning benchmarks have been available on the Web since before these models were trained, this possibility cannot be ruled out by using existing benchmarks such as GSM8K or MATH. Moreover, while these benchmarks are useful for obtaining aggregate estimates of reasoning capabilities, their diversity also precludes a more fine-grained understanding of how specific mathematical skills might evolve or interact. For instance, while “grade-school math” (e.g., in GSM8K) consists of an extremely broad set of specific skills, existing datasets aggregate all of these into a single final accuracy number. This is useful for comparing different models, but less informative about what exactly models learn or how they might differ.
In this work, we develop a synthetic dataset and design evaluations to address these questions, and elicit a number of insights into how mathematical skills develop during LLM pre- and post-training. Concretely, we introduce MathCAMPS: a fine-grained dataset created automatically from the Common Core (CC) mathematics curriculum, including 49 different skills (CC “standards”) from grades K-8. The CC curriculum is adopted by thousands of K-12 schools and includes grade-wise standards that indicate specific skills students must be proficient at by grade. By constructing MathCAMPS in direct relation to the CC, our benchmark enables rich set of analyses of mathematical proficiency in language models, allowing direct parallels to abilities that human students are also evaluated on.
We encode the family of problems associated with each CC standard as a grammar, allowing us to sample arbitrarily many symbolic problems, which our pipeline then converts into natural language word problems using GPT-4o (OpenAI et al., 2024). To ensure that the word problems are faithful to the symbolically generated structures, we introduce a method for cycle consistency that allows us to ensure data quality in a fully automatic manner.
@@
Our fine-grained dataset then allows us to study the training dynamics of several open-weight language models for which training checkpoints have been made available. We report on a series of novel analyses of the learning dynamics of mathematical reasoning skills that our methodology enables. Our contributions are:
  1. 1.
    We construct MathCAMPS, a dataset of 4900 new problems stratified into fine-grained capabilities defined by the Mathematics Common Core Standards for K-8 grades. We also release our pipeline to generate arbitrarily many more.
  2. 2.
    We use OLMo2, Amber, and Pythia (families of open-weight models) to analyze how specific mathematical skills evolve during pre-training. We explore alignment between LLM and human learning, skill evolution patterns, and the robustness of reasoning abilities.
  3. 3.
    We use OLMo2 and Amber models to explore how instruction tuning, a widely used post-training method, affects specific mathematical abilities. We find that the specific instruction tuning methodology used in each model varies widely into how broadly beneficial they are to reasoning, as well as which specific mathematical skills they affect the most — both positively and negatively.
@@ 2 Related Work
Our work closely relates to (i) current benchmarks of mathematical reasoning in LLMs, (ii) benchmarks constructed using LLMs, (iii) behavioral testing and applications in NLP, and finally (iv) open-weight LLMs that provide training checkpoints.
@@
MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021) have been two leading benchmarks for the evaluation of mathematical reasoning in LLMs. Both datasets consist entirely of human-authored problems — a process that is expensive to reproduce — and as a result, neither benchmarks were updated since their initial releases. Given that LLMs are trained on Web data, it is unclear whether they might have been trained on the test problems of these benchmarks (Bubeck et al., 2023) – either directly or from other sources (e.g., all problems in MATH come from past public competitions). In fact, GSM1K (Zhang et al., 2024), a new dataset that independently attempted to reproduce the data distribution of GSM8K, has found reduced performance on several models, suggesting the possibility of test set contamination. Moreover, these datasets evaluate a very wide range of abilities, without distinction. The GHOSTS dataset provided the first fine-grained evaluation of undergraduate-level mathematical skills in LLMs (Frieder et al., 2023); but given that the reasoning evaluation was performed manually by expert annotators, this unfortunately does not scale to evaluating a large number of training checkpoints. In contrast, the evaluation in MathCAMPS is both fine-grained and fully automated, albeit at the grade school level.
@@
As collecting data from human annotators at scale is expensive (especially in domains requiring expertise), prior work has relied on LLMs to aid the generation of large-scale benchmarks (Hartvigsen et al., 2022). BigToM (Gandhi et al., 2023), a benchmark of social reasoning in LLMs, applied the idea of symbolically scaffolding questions for the LLM to realize in natural language, an approach that we transport to mathematics. Dyval (Zhu et al., 2024) proposed a method for generating reasoning problems for LLMs based on a DAG representing the computation. While Dyval contains two mathematical tasks (arithmetic and solving linear equations), MathCAMPS takes this idea further for mathematical reasoning, spanning 44 skills directly grounded on a human curriculum. Other synthetic evaluations focused on mathematical skills include GSMore (Hong et al., 2024) and the concurrent work on GSM-Symbolic (Mirzadeh et al., 2024). Both these works focus on evaluating the robustness of LLMs by perturbing existing problems from an existing dataset, GSM8k, whereas in MathCAMPS we synthesize problems from scratch, grounded on a human curriculum (Hong et al. (2024) also proposes perturbations to coding problems, which we do not focus on here).
@@
Our goal to provide a fine-grained evaluation of mathematical reasoning has parallels with behavioral testing — the idea of testing software systems on specific features, as opposed to just their overall adequacy (Ribeiro et al., 2020). In particular, CheckList (Ribeiro et al., 2020) allowed testing machine translation models for fine-grained failure modes. Dynaboard (Ma et al., 2021) proposed an NLP leaderboard where users can adapt to their own needs by choosing the utility of different metrics; our dataset enables a similar user-customizable comparison between models for mathematical reasoning.
Open-Weight Language Models
Our analysis of the learning dynamics of is done on top of models for which training checkpoints are made publicly available. These include Pythia (Biderman et al., 2023), OLMo (OLMo et al., 2024), and LLM360 (Liu et al., 2023). Moreover, OLMo and LLM360 also have post-trained (instruction-tuned) versions, allowing us to also understand the impact of this step in the LLM training pipeline.
@@ 3.1 The Mathematics Common Core
We take 44 standards spanning grades K through 8 to compose MathCAMPS, focusing on standards that are amenable to automatic problem generation with a final answer in text form. The complete CC curriculum has 229 standards across grades K through 8, bringing our coverage to 19.2% of the curriculum for these grades. Notably, we currently do not cover standards focusing on conceptual understanding (e.g., 3.OA.D.9 – “Identify arithmetic patterns […], and explain them using properties of operations.”), or standards that emphasize visual reasoning (e.g., 6.G.A.4 – “Represent three-dimensional figures using nets made up of rectangles and triangles, and use the nets to find the surface area of these figures.”). All 44 standards covered in MathCAMPS are listed in Appendix C.
@@
We represent CC standards as non-terminals in an attribute grammar (Heine & Kuteva, 2007) — a rich formalism that can encode semantic, context-sensitive rules. Attribute grammars can encode syntax much like a context-free grammar, but also allow us to embed information processing (e.g., setting and testing conditions on attributes, such as bounds on constants) in the production rules. We map each standard $s$ to a non-terminal $P_{s}$, such that all strings produced by expanding $P_{s}$ using production rules are valid symbolic representations of a problem pertaining to standard $i$. Figure 1 shows a (simplified) grammar for the standard 1.OA.A.1 – “Use addition and subtraction within 20 to solve word problems involving situations of adding to, taking from, putting together”. Here, a word problem, generated by the Problem non-terminal, consists of a sequence of declarative statements expressing equations between expressions. For this standard, an expression consists of addition, subtraction, variables, and constants. After these declarations, the problem ends with a question — an expression representing the value that the problem asks for. Concretely, our grammar is implemented in Python: each non-terminal becomes a stochastic function that samples and applies a production rule, recursively expanding non-terminals that it produces. In the grammar in Figure 1 (A), sampling a Problem generates a structure such as the one shown in Figure 1 (B).
@@
When sampling problems, there is no a priori guarantee that all generated statements are necessary to answer the question. To avoid such statements, we remove them by applying a simple graph reachability algorithm on a dependency graph between statements, removing statements that the answer does not depend on. This enforces the constraint of only having useful statements in problems. Besides this constraint, which we always enforce, each standard can apply specific constraints. The standard 1.OA.A.1 has an example of such constraint: it requires that students only be asked to use “addition and subtraction within 2020.” To be faithful to this standard, we must validate that no intermediate values used in the solution exceed $20$. To encode this and other constraints across the curriculum, we implement a suite of $6$ parameterized filters (detailed in Appendix D) that are selectively applied depending on the standard’s specification. Applying rejection sampling from the grammar using the standard’s filters gives a procedure for generating valid symbolic problems. For all standards that can be formulated as solving a system of linear equations, we use SymPy (Meurer et al., 2017) to obtain final answers. For other cases, we use two simple custom procedures (to list the factors of numbers and to compare values).
@@ 3.2 From symbolic to word problems
Unlike generating problem stories from a fixed set of templates, using a language model for generating natural language problems gives us fluid, diverse language. Unfortunately, we also lose any guarantee that the generated word problem represents the original symbolic structure faithfully. To mitigate this issue, we also introduce a cycle consistency method that we have found to drastically improve problem quality. Precisely, we use the same few-shot examples we crafted for each standard in reverse (i.e., with the natural language problem coming first, followed by the symbolic structure) to have GPT-4 translate the word problem it wrote into a symbolic structure. In this step, the model is not given the original structure. We then parse and apply the appropriate solver to the generated symbolic problem; we consider the generation cycle-consistent if the answers to the original and recovered problems are the same (illustrated in Figure 1). We then discard problems that fail this test. A more in-depth analysis of the efficacy of cycle-consistency can be found in Appendix F.